# Optimal Transmission Power Scheduling for Networked Control System under DoS Attack

### v1

@@ V Simulation
In this section, we provide numerical examples to illustrate our results. Consider a second-order system with system dynamics $A={\rm diag}\{1.3,-1.1\}$, $B=[0.1~{}0.1]^{\top}$, process noise covariance $W=0.001$. The weighting matrices in the LQG function (5) are chosen as $Q=R={\rm diag}\{1,1\}$. The discount factor is chosen as $\gamma=0.9$. The covariance of the initial value is chosen as $X_{0}={\rm diag}\{0.01,0.01\}$.
Figure 2: From top to button: attack energy and transmission power; transmission success index under greedy scheduler (21).
Figure 3: From top to button: system state; control signal under the greedy scheduler (21).
@@
Figure 5: The total cost achieved by the greedy scheduler (21), the approximation-based greedy scheduler (22). The theoretical and empirical upper bounds of the total cost achieved by the designed schedulers. The theoretical lower bound of the total cost achieved by the best constant-power scheduler.
The communication channel parameters are chosen as $\sigma^{2}=1$, $\alpha=3$, see [17]. The attack energy is chosen as a uniformly distributed random variable between $[0,1]$.
@@
We choose the time horizon as $T=100$ and the tradeoff multiplier as $\lambda=1$. The top subfigure of Fig. 2 depicts the energy of the attack and the transmission power determined by scheduler (21) during $k\in\mathbb{N}_{[0,T]}$. The button subfigure of Fig. 2 depicts the transmission success index during $k\in\mathbb{N}_{[0,T]}$. It can be observed from Fig. 3 that transmission success is affected by both the transmission power and attack energy. When addressing the resource utilization efficiency under attacks, a higher attack energy generally leads a higher transmission power. Fig. 3 depicts the system state trajectory, its estimate, and control signal. It can be observed from Fig. 3 that the remote estimator effectively tracks the system dynamics. Fig. 4 depicts the mean square error under different average transmission power achieved by the greedy schedulers (21) and (22) and the scheduler with constant transmission power $p_{k}=3$ for $k\in\mathbb{N}_{[0,T]}$. The mean square error is measured by $\frac{1}{T}\sum_{k=0}^{T}e_{k}^{\top}A^{\top}\Sigma Ae_{k}$ and the average transmission power is measured by $\frac{1}{T}\sum_{k=0}^{T}p_{k}$. It can be observed that using the same average transmission power, greedy schedulers (21) and (22) achieves smaller mean square error than constant-power schedulers.
@@
Fig. 5 depicts the total regulation and transmission costs under different multipliers under the greedy scheduler (21), the approximation-based greedy scheduler (22) assuming attack energy is unknown, and constant-power schedulers. We choose tradeoff multiples $\lambda\in[0.01,1]$. Monte Carlo simulation runs $20000$ trials. As in (17), only last-second terms depend on the designed control and scheduling laws. We choose $\sum_{k=0}^{T}\big{(}\gamma^{k}e_{k}^{\top}\Sigma e_{k}+\lambda\gamma^{k}p_{k}\big{)}$ as the empirical total regulation and transmission cost. For the empirical results, shaded areas represent $\pm$ one standard deviation over 20000 trials. Accordingly, the theoretical upper bound of the total cost, i.e., the cost achieved by a transmission scheduler with maximum constant power, is $\frac{\gamma{\rm{tr}}(\Theta W)+\lambda p_{\max}}{1-\gamma}+{\rm{tr}}(\Theta X_{0})$. The empirical upper bound is calculated by implementing the constant-power scheduler $p_{k}=3$ for all $k$. It can be observed from Fig. 5 that the theoretical upper bound of the total cost approaches the empirical one. As a comparison, we also depict the lower bound of the total cost by a constant-power scheduler, i.e., the best performance it can achieve, which is given by
$\min_{q\in\mathcal{Q}}\bigg{\{}\frac{\gamma{\rm{tr}}(\tilde{\Theta}W)}{1-\gamma}+{\rm{tr}}(\tilde{\Theta}X_{0})+\frac{\lambda\mathbf{E}_{a\sim\mathcal{D}_{a}}[p(q,a)]}{1-\gamma}\bigg{\}},$
${\rm s.t.}\ \gamma qA^{\top}\tilde{\Theta}A+\Sigma=\tilde{\Theta}.$
Moreover, it can be observed that greedy schedulers (21) and (22) both outperform the schedulers with constant power, which includes the one that achieves the minimum cost among all constant-power schedulers.
@@ VI Conclusion
In this article, we have studied the optimal co-design of control law and transmission power scheduler that minimizes the regulation and transmission costs for networked control systems under DoS attacks. Given the acknowledgment signal from the remote controller, the information structure between the controller and the power scheduler is nested. Therefore, we showed that the original co-design can be decomposed into the optimal control design, yielding a certainty equivalence controller, and the optimal power scheduling design, which is tracked by dynamic programming approaches. Expressions of the optimal power scheduling were provided in both finite- and infinite-horizon cases. To ease the computational complexity in finite-horizon dynamic programming, an alternative greedy scheduler was developed for implementation. Additionally, in the infinite-horizon case, we provided the upper bound of the total regulation and transmission cost under the proposed scheduler.

### v2

@@ V Simulation
In this section, we provide numerical examples to illustrate our results. Consider a second-order system with system dynamics $A={\rm diag}\{1.3,-1.1\}$, $B=[0.1~{}0.1]^{\top}$, process noise covariance $W=\mathrm{diag}\{0.001,0.001\}$. The weighting matrices in the LQG function (5) are chosen as $Q=R={\rm diag}\{1,1\}$. The discount factor is chosen as $\gamma=0.9$. The covariance of the initial value is chosen as $X_{0}={\rm diag}\{0.01,0.01\}$. The communication channel parameters are chosen as $\sigma^{2}=1$, $\alpha=3$, see [24]. The attack energy is chosen as a uniformly distributed random variable between $[0,1]$, i.e., $a_{k}\sim U[0,1]$. We choose the time horizon as $T=100$ and the tradeoff multiplier as $\lambda=1$.
Figure 2: From top to bottom: attack energy and transmission power; transmission success index under greedy scheduler (20).
Figure 3: From top to bottom: system state; control signal under the greedy scheduler (20).
@@
Figure 5: The total cost achieved by the greedy scheduler (20), the approximation-based greedy scheduler (21). The theoretical and empirical upper bounds (34) of the total cost. The theoretical and empirical cost achieved by constant-power scheduler with $p_{k}=3$.
Figure 6: The empirical average cost achieved by the greedy scheduler (20) and the corresponding theoretical upper bounds (34) with different attack energy distributions.
@@
The top subfigure of Fig. 2 depicts the attack energy and the transmission power determined by scheduler (20) for $k\in\mathbb{N}_{[0,T]}$. The bottom subfigure of Fig. 2 depicts the transmission success index for $k\in\mathbb{N}_{[0,T]}$. Fig. 2 shows that transmission success is affected by both the transmission power and attack energy. Additionally, a higher attack energy generally leads to higher transmission power. Fig. 3 depicts the system state trajectory, its estimate, and control signal. Fig. 3 shows that the remote estimator effectively tracks the system dynamics. Fig. 4 depicts the mean square error achieved by the greedy schedulers (20) and (21) and the constant-power schedulers, with the average transmission power $p=[0.4,1.8]$. The mean square error is measured by $\frac{1}{T}\sum_{k=0}^{T}e_{k}^{\top}A^{\top}\Sigma Ae_{k}$ and the average transmission power is measured by $\frac{1}{T}\sum_{k=0}^{T}p_{k}$. Fig. 4 shows that for the same average transmission power, the performance of greedy schedulers (20) approaches that of approximation-based scheduler (21), which aligns with Remark 2. Furthermore, they outperform the constant-power schedulers in achieving smaller mean square error.
@@
Fig. 5 depicts the total regulation and transmission costs (16) with different multipliers under the greedy scheduler (20), the approximation-based greedy scheduler (21), and constant-power schedulers. We choose tradeoff multiples $\lambda\in[0.01,1]$. Monte Carlo simulation runs $20000$ trials. We choose $\sum_{k=0}^{T}\big{(}\gamma^{k}e_{k}^{\top}\Sigma e_{k}+\lambda\gamma^{k}p_{k}\big{)}$ as the empirical total regulation and transmission cost. For the empirical results, shaded areas represent $\pm$ one standard deviation over 20000 trials. A small shaded area means that the designed scheduling policy delivers steady performance. The theoretical upper bound of the cost, i.e., the cost achieved by the optimal constant-power scheduling $\tilde{\pi}$, is $\frac{\gamma{\rm{tr}}(\tilde{\Theta}W)+\lambda\tilde{p}}{1-\gamma}+{\rm{tr}}(\tilde{\Theta}X_{0})$, where $\tilde{p}$ and $\tilde{\Theta}$ are defined in (35). The theoretical cost achieved by a transmission scheduler using constant power $p_{k}=p_{\max}=3$ is $\frac{\gamma{\rm{tr}}(\hat{\Theta}W)+\lambda p_{\max}}{1-\gamma}+{\rm{tr}}(\hat{\Theta}X_{0})$, where $\hat{\Theta}$ satisfies $\gamma\hat{q}A^{\top}\hat{\Theta}A+\Sigma=\hat{\Theta}$. Fig. 5 shows that the theoretical costs $\hat{\Psi}(f^{\ast},\tilde{\pi})$ and $\hat{\Psi}(f^{\ast},\hat{\pi})$ both match their corresponding empirical costs, which illustrates the effectiveness of theoretical bound calculation. Additionally, it shows that the greedy schedulers (20) and (21) both outperform arbitrary constant-power schedulers, including the optimal one. Furthermore, the greedy scheduler (20) achieves the performance comparable to that of the approximation-based greedy scheduler (21), echoing the statement in Remark 2.
Fig. 6 depicts the total regulation and transmission costs achieved by greedy scheduler (20) and its theoretical upper bounds (34) under different attack distributions. We choose two uniform distributions $U_{1}[0,1]$ and $U_{2}[0,2]$; two Poisson distributions $\mathcal{D}_{p}^{\lambda}$ with the p.d.f. of $f(k,\lambda)=\mathrm{Pr}(X=k)=\frac{\lambda^{k}e^{-\lambda}}{k!}$, where $\lambda\in\{1/2,1\}$. Define the truncated normal distribution with p.d.f $f(x,\mu,\sigma,a,b)=\frac{1}{\sigma}\frac{\varphi(\frac{x-\mu}{\sigma})}{\Phi(\frac{b-\mu}{\sigma})-\Phi(\frac{a-\mu}{\sigma})}$ for $x\in[a,b]$ and $f=0$ otherwise, where $\varphi$ is the p.d.f of the standard normal distribution and $\Phi(x)=\frac{1}{2}(1+{\mathrm{erf}}(x/\sqrt{2}))$. We choose two truncated normal distribution $\mathcal{N}_{t}^{1}$ and $\mathcal{N}_{t}^{2}$ with $f(x,1/2,1/12,0,1)$ and $f(x,1,1/3,0,2)$. Note that the distributions $U_{1}$, $\mathcal{D}_{p}^{1/2}$ and $\mathcal{N}_{t}^{1}$ have the same mean value, and distributions $U_{2}$, $\mathcal{D}_{p}^{1}$ and $\mathcal{N}_{t}^{2}$ have the same mean value. Monte Carlo runs 20000 trials. For clarity, we only plot the mean value of the empirical costs and omit their standard deviations. Fig. 6 shows that, given the same average attack energy, the greedy scheduler achieves the same performance across various attack energy distributions. This observation aligns with Remark 2. Moreover, under the same scheduling policy, the total cost increases as the attack energy increases. Additionally, theoretical upper bounds successfully constrain the empirical costs under different attack energy distributions.
@@ VI Conclusion
In this article, we have studied the optimal co-design of control law and transmission power scheduler that minimizes the regulation and transmission costs for networked control systems under DoS attacks. Given the acknowledgment signal from the remote controller and same knowledge about the attack energy, the information structure between the controller and the power scheduler is nested. Then, we showed that the original co-design can be decomposed into the optimal control design, yielding a certainty equivalence controller, and the optimal power scheduling design, which is tracked by dynamic programming approaches. Expressions of the optimal power scheduling were provided in both finite- and infinite-horizon cases. To ease the computational complexity in finite-horizon dynamic programming, an alternative greedy scheduler was developed for implementation. Additionally, in the infinite-horizon case, we provided the upper bound of the total regulation and transmission cost under the proposed scheduler. Nevertheless, the proposed design relies on specific assumptions about the nature of DoS attacks, such as magnitude or distribution, which may limit its applicability in more diverse or unpredictable attack scenarios. As a result, its effectiveness may be limited when facing more sophisticated or unpredictable attack patterns. To address this limitation, future research will focus on developing adaptive power scheduling mechanisms that can dynamically adjust in response to real-time detection of evolving or unknown DoS attacks.